# Spatially covariant gravity with two degrees of freedom in the presence of an auxiliary scalar field: Hamiltonian analysis

### v1

@@ Appendix B Constraint algebra
In this appendix, we present the explicit expressions for the Poisson brackets.
@@
$\left[\mathcal{C}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $2B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}-\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta N(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}),$ (143)
@@ Appendix C Adjoint matrix form of null eigenvector
If we wish to find the null eigenvector of the degenerate matrix $\mathcal{B}^{ab}$, one method is to calculate the non-zero rows (or columns) of its corresponding adjoint matrix $\mathcal{B}_{ab}^{*}$, and to multiply the corresponding elements by $(-1)^{a+b}$ to build the null eigenvector. For simplicity, we assume that $\mathcal{B}^{ab}$ is a simple $\times 3$ 3 matrix and the rows and columns correspond to constraints $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$.

### v2

@@ Appendix B Constraint algebra
In this appendix, we show the explicit expression for the Poisson brackets.
@@
$\left[\mathcal{C}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $2B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}-\int\mathrm{d}^{3}z\frac{\delta S}{\delta N(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}),$ (143)
@@ Appendix C Adjoint matrix form of null eigenvector
If we wish to find the null eigenvector of the degenerate matrix $\mathcal{B}^{ab}$, one method is to calculate the non-zero rows (or columns) of its corresponding adjoint matrix $\mathcal{B}_{ab}^{*}$, and to multiply the corresponding elements by $(-1)^{a+b}$ to build the null eigenvector. For simplicity, we assume that $\mathcal{B}^{ab}$ is a simple $3\times 3$ matrix and the rows and columns correspond to constraints $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$.